Entire Session

Session: Feb 2014 (session 1)
Wavelengths: 18 5 6 1.3 cm


To give feedback on this entire session, click here

Entire Session Feedback

Onsala (Jun Yang at Mon Mar 03 12:11:54 2014) Success: RFI
There are some strong spikes at frequency close to 1667MH in the LCP light path of our new L-band receiver. Once the session is end, our engineers will start to fix the minor problem.

Metsahovi (Minttu Uunila at Wed Mar 12 11:30:05 2014) Minor failures: TSYS
Negative Tsys values throughout the session.